OpenStudy (anonymous):

How do I calculate (-5p^4 z^6 u)^3

OpenStudy (anonymous):

Power each terms!

OpenStudy (anonymous):

I don't know how to power the terms

OpenStudy (anonymous):

like 2² = 2* 2 = ...?

OpenStudy (anonymous):

Makes sense

OpenStudy (anonymous):

( -5p⁴)³ = ( -5)³ * (p⁴)³ = .....